# Learning to Control Unknown Strongly Monotone Games

### v1

@@ 6 Conclusion
From a technical point of view, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$. Our algorithm essentially controls the unknown game such that its NE becomes desirable from an optimization point of view. Since convergence to NE is relatively fast, this has the potential to accelerate general cooperative multi-agent optimization [62].
Controlling unknown games with bandit feedback is a new online learning paradigm. The game to be controlled can be thought of as a non-stationary but highly structured bandit. Viewing game control as a bandit problem can result in both new exciting research questions, and new powerful game control techniques.
Generalizing our game control setting beyond linear constraints is an exciting new research avenue that can help tackle a wider class of problems.

### v2

@@ 6 Conclusion
Technically, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$. Our algorithm essentially controls the unknown game such that its NE becomes desirable from an optimization point of view. Since convergence to NE is relatively fast, this has the potential to accelerate general cooperative multi-agent optimization [60].
Controlling unknown games with bandit feedback is a new online learning paradigm. The game to be controlled is a non-stationary but highly structured bandit. Viewing game control as a bandit problem can result in new exciting research questions, and new powerful game control techniques. Generalizing our game control setting beyond linear constraints can help tackle a wider class of problems.